OpenStudy (anonymous):

Solve the equation for the indicated variable. 3x + 7y = 2, solve for y

OpenStudy (chealyn98):

3x + 7y = 2 -3. -3 ------------------ 7y = -1 ---- ----- 7 7 Then just divided

OpenStudy (anonymous):

Ok thank you so much!!

OpenStudy (chealyn98):

Your welcome